OpenStudy (anonymous):

Simplify -5+i over 2i

OpenStudy (solomonzelman):

\[\frac{(-i+5)\color{red} {\times i}}{2i\color{red} {\times i}}\]\[\frac{1+5i}{-2}~~~~~->~~~~~-\frac{1+5i}{2}\]

OpenStudy (anonymous):

Okay, so you multiply both the top and bottom by I, and that gives you 1+5i over -2?

OpenStudy (solomonzelman):

that's all.

OpenStudy (anonymous):

So is the final thing 1+5i over 2? and the whole thing is negative?

OpenStudy (solomonzelman):

yes, it's \[\huge\color{blue}{-\frac{1+5i}{2}}\]

OpenStudy (solomonzelman):

Oh, ty for the metal!

OpenStudy (anonymous):

Can you help me with another one?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

Please one question per thread in the future. I will help you with this one though.